what is the debroglie wavelength of an antineutrino with a kinetic energy of 1.8 MeV? take the mass of the antineutrino to be 4.99E-37 kg.
Answer #1

Given :K=1.8MeV=1.8*1.6E-13 J=2.88E-13 J

M=4.99E-37 kg

Momentum:p=√2MK=√2*4.99E-37*2.88E-13

p=5.36E-25 kg-m/s

De-broglie wavelength: Y=h/p

Y=(6.626E-34)/(5.361E-25)

Y=(1.236E-9)m=1.236nm